A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.4 MB. Yesterday, there were 1010 downloads of the song, for a total download size of 3424 MB. How many downloads of the high-quality version were there?
*Really need help on this problem. Tried it several times and could not get the answer. Please help. Thanks! :)

Answer by Theo(13342)

let x = number of standard quality downloads at 2.9 MB.
let y = number of high quality downloads at 4.4 MB.

x + y = 1010
2.9x + 4.4y = 3424

these 2 equations need to be solved simultaneously either by graphing or by substitution or by elimination or by matrix methods which are similar to elimination.

we'll do elimination.

mjultiply both sides of the first equation by 2.9 to get:

2.9x + 2.9y = 2929
2.9x + 4.4y = 3424

subtract the first equation from the second eqution to get:

1.5y = 495

divide both sides by 1.5 to get y = 330

since x + y = 1010, this means that x has to be equal to 680 because 680 + 330 = 1010.

since you used the first equation to solve for x, then use the second equation to confirm your solutions are correct by replacing x with 680 and y with 330 in that equation to get:

2.9x + 4.4y = 3424 becomes 2.9*680 + 4.4*330 = 3424 which becomes 1972 + 1452 = 3424 which becomes 3424 = 3424, confirming your solution is correct.

your solution is:

number of high quality downloads was 330.